What's the future value of an investment of $1 a year for each of 4 years, at the end of the last year? Suppose the interest rate is 8%.

Multiple Choice

  • $5.87

  • $4.51

  • $3.25

  • $5.64
Answer #1

Future value of the annuity formula:

\small \textup{FVA} = A\left [ \frac{(1+i)^{n}-1}{i} \right ]

Where,
FVA = Future Value of Annuity
A = Annuity or yearly deposit
i = rate of interest in decimal form (i.e 8% = 0.08)
n = number of years

Therefore,

\small \textup{FVA} = 1\left [ \frac{(1+0.08)^{4}-1}{0.08} \right ]

\small = 1\left [ \frac{(1.08)^{4}-1}{0.08} \right ]

\small = 1\left [ \frac{1.36049-1}{0.08} \right ]

\small = 1\left [ \frac{0.36049}{0.08} \right ]

\small = \$4.51